shift summary for owl shift, 7/14

Jim Hirschauer (jhirsch@cuhep.Colorado.EDU)
Wed, 14 Jul 1999 06:05:28 -0600 (MDT)

Shift Summary for owl shift, 14/Jul/99

On Shift: Peter Shanahan, Jim Hirschauer

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 2

Accounting of lost time: No beam.

Runs ended this shift:

Run 13899 - KTEV_PED

Tapes brought to FCC this shift:

Accesses during shift: Continuous accesses by EB, RK, SET, JFH
from 00:00 to 4:45 to fix HA shaper/splitter

Overview of shift: No beam due to MI Quad swap/other MI problems

Description of Shift:

run spill time comment
00:30 - 04:45 EB, RK, SET, JFH on access to work on HA PS
problem see ~da_ktev/doc/RR90_PROBLEMS.TXT for
the whole story. Bottom line: HA
shaper/splitters are bypassed. Only analog sum
is ADC'd now.

4:20 SET: changing the base on BA 77 doesn't fix
problem, conclusion: probably bad PMT, since
changing PMT is risky it will not be attempted.
Changing the base one more time may be

5:00 When keys are returned to the key tree, "Permit
Loop A" light does not come on -- problem solved
by methodical jiggling of keys.

5:30 Beam stop is disabled, magnets are set for muon
mode. We are only awaiting beam.

5:54:29 "Beam fallout in MI. Experts investigating. No
estimate for beam."

Follow up Notes on Previous Shift

Items to Follow up on Next Shift

o Take muon run (magnets are now setup for muon running)
o keep collecting BEAM_832 data (NM4AN is "POS". There seems to be some
confusion about this in the log book. Check with run manager.)
o Page run manager when beam arrives.